# DiT4SR: Taming Diffusion Transformer for Real-World Image Super-Resolution

### v1

@@ 2 Related Work
With the success of large-scale T2I models such as Stable Diffusion (SD) [37], researchers [34, 61, 45, 19, 39, 15, 43, 40] have shifted their attention to use the generative priors embedded in these powerful pre-trained models to handle the Real-ISR problem. Built upon the first-generation SD [37], which is based on a UNet architecture, StableSR [47] and DiffBIR [29] inject the LR information as conditions via ControlNet [64] or ControlNet-like manners. PASD [59] and SeeSR [54] further incorporate high-level semantic information to guide the diffusion process. Employing the more advanced SDXL [33], SUPIR investigates the effects of scaling up in Real-ISR. There are also some other methods [50, 62, 24, 57, 7, 53, 41, 30, 31] committed to a more efficient diffusion process. With the development of diffusion transformer (DiT), DreamClear [2] propose a DiT-based image restoration model, but still adopts the ControlNet to inject LR information, which prevents it from fully leveraging the advantage of DiT. Differently, our DiT4SR is one of the pioneering works to tame the large-scale diffusion transformer model for Real-ISR, while discarding the ControlNet-like approach.

### v2

@@ 2 Related Work
With the success of large-scale T2I models such as Stable Diffusion (SD) [41], researchers [38, 65, 49, 21, 43, 16, 47, 44, 26] have shifted their attention to use the generative priors embedded in these powerful pre-trained models to handle the Real-ISR problem. Built upon the first-generation SD [41], which is based on a UNet architecture, StableSR [51] and DiffBIR [33] inject the LR information as conditions via ControlNet [68] or ControlNet-like manners. PASD [63] and SeeSR [58] further incorporate high-level semantic information to guide the diffusion process. Employing the more advanced SDXL [37], SUPIR investigates the effects of scaling up in Real-ISR. There are also some other methods [54, 66, 27, 61, 7, 57, 45, 34, 35] committed to a more efficient diffusion process. With the development of diffusion transformer (DiT), DiT-SR [14] trains DiT-based SR models from scratch. DreamClear [2] proposes a DiT-based image restoration model, but still adopts the ControlNet to inject LR information, which prevents it from fully leveraging the advantage of DiT. From the perspective of network architecture, our DiT4SR is one of the pioneering works to tame the large-scale diffusion transformer model for Real-ISR, while discarding the ControlNet-like approach. There are two concurrent methods [28, 19] that explore one-step DiT-based SR models in terms of flow trajectory.